Ida Sidha Karya Company is a family-owned company located in the village of Gianyar on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold for S794. Selected data for the companys operations last year follow Units in beginning inventory Units produced Units sold Units in ending inventory Variable costs per unit: 25,000 22,000 3,000 Direct materials Direct labor Variable manufacturing overhead Variable selling and administrative $ 170 S 430 S 60 S 19 Fixed costs: Fixed manufacturing overhead Fixed selling and administrative S 840,000 610,000 Required 1. Assume that the company uses absorption costing. Compute the unit product cost for one gamelan.(Round your intermediate calculations and final answer to nearest whole dollars.) t cost 2. Assume that the company uses variable costing. Compute the unit product cost for one gamelan. it product cost
Answer #1

Answer

  • All working forms part of the answer
  • Requirement 1

Product cost under absorption costings includes direct costs and manufacturing overheads, including fixed manufacturing overheads.

A

Fixed manufacturing Overhead

$              840,000

B

Units produced

                    25,000

C= A / B

Fixed manufacturing cost per unit

$                         34

Variable manufacturing costs:

D

Direct Material

$                      170

E

Direct Labor

$                      430

F

Variable manufacturing overhead

$                         60

G = C+D+E+F

Unit product Cost

$                      694

  • Requirement 2

Under Variable costing, Unit Product cost only includes variable manufacturing costs.

Variable manufacturing costs:

A

Direct Material

$                    170

B

Direct Labor

$                    430

C

Variable manufacturing overhead

$                      60

D = A+B+C

Unit product Cost

$                    660

  • Note: Selling & administrative expense [Fixed and variable] do not forms part of product cost. These are Period Cost.